OpenStudy (anonymous):

Solve and determine if a solution is extraneous. Can someone please explain.Thanks! In progress...

OpenStudy (anonymous):

\[\sqrt{x+16}+\sqrt{x=8}\]

OpenStudy (whpalmer4):

I think that is probably \[\sqrt{x+16} + \sqrt{x} = 8\]Correct?

OpenStudy (anonymous):

yes correct sorry

OpenStudy (jhannybean):

so solve for x we could square each side.

OpenStudy (anonymous):

yes that i kind of understand after squared then what?

OpenStudy (anonymous):

\[\left( \sqrt{x+16}+\sqrt{x=8}\right)^2\]

OpenStudy (loser66):

I am not sure whether I am right or wrong, just try. watch me, please let x = a^2 . I try to form the form of |a +4i| = sqrt (a^2 +4^2| and the second one is sqrt(a^2) = |a|

OpenStudy (loser66):

then |a+4i|+|a|= 8 solve for a, plug back to solve for x

OpenStudy (jhannybean):

\[\large \sqrt{x+16} + \sqrt{x} = 8\] square both sides \[\large (\sqrt{x+16} + \sqrt{x})^2 = (8)^2\] expand \[\large (\sqrt{x+16} +\sqrt{x})(\sqrt{x+16}+\sqrt{x}) = 64\]

OpenStudy (dan815):